OpenStudy (anonymous):

does anyone understand this finding the real solution? the 3 and 2 are exponets ???x^3/^2 -27=0

OpenStudy (shadowfiend):

Hm. What do you mean? The real solution of each is the appropriate root. e.g.: $$ x^2 - 27 = 0 $$ $$ x^2 = 27 $$ $$ x = \pm \sqrt{27} $$

OpenStudy (shadowfiend):

Ah, sorry, misunderstood. You have \(x^{\frac{3}{2}}\) In that case the process is similar, but you have to remember that \(\sqrt{x}\) is the same as \(x^2\), so \(x^{\frac{3}{2}}\) is really the same as \(\sqrt{x^3}\). So, you end up with: $$ \sqrt{x^3} = 27 $$ You can reverse the square root by squaring and the cube by taking the cubed root.

OpenStudy (shadowfiend):

I meant \(\sqrt{x}\) is the same as \(x^{\frac{1}{2}}\).

OpenStudy (anonymous):

Thanks so much:)